OpenStudy (anonymous):

Can anyone help me with this equation? Use Quadratic Formula -2x2 - 5x + 5 = 0

OpenStudy (ybarrap):

$$ x=\cfrac{-b\pm\sqrt{b^2-4ac}}{2a} $$ Where $$ a=-2\\ b=-5\\ c=5 $$ This is the quadratic formula. There will be two answers, due to the \(\pm\) part.

OpenStudy (ybarrap):

$$ ax^2+bx+c=c $$ Note that your equation is in this form, with the values for \(a,b\) and \(c\) above.

OpenStudy (ybarrap):

$$ x_1=x=\cfrac{-b+\sqrt{b^2-4ac}}{2a} $$ Will be one solution. The other will be, $$ x_2=x=\cfrac{-b-\sqrt{b^2-4ac}}{2a} $$ That it!

OpenStudy (anonymous):

thank you :)

OpenStudy (ybarrap):

$$ ax^2+bx+c=0 $$ Note that your equation is in this form, with the values for \(a,b\) and \(c\) above.

OpenStudy (ybarrap):

That should have been a zero there at the end. NP